OpenStudy (anonymous):

Find all solutions for 3sin^2 θ +16 sin θ -35=0. (Use 'k' as any integer)

OpenStudy (anonymous):

Replace sin(theta) with x \[3x ^{2}+16x-35=0\] \[(3x - 5)(x + 7)\=0\]

OpenStudy (anonymous):

Now replace back: \[(3\sin \theta-5)(\sin \theta+7)=0\] So either \[3\sin \theta-5=0\] or \[\sin \theta+7=0\] \[\sin \theta \neq-7\] so \[3\sin \theta=5\] \[\sin \theta=\frac{ 5 }{ 3 }\] but this can't be either since \[-1\le \sin \theta \le1\] It seems as though we have "No Solution"